40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 118 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  11041   Tue Feb 17 00:24:47 2015 rana, jenneUpdateLSCALS Fool filter updated for more cancellation

Today we measured the TFs again and then updated the filter in the POY -> ALS FF path so as to get 10x better cancellation.

The cancellation went from ~10 dB to ~30 dB. This seems good enough. The new filter 'Comp1' is just constructed by eye. We then had to tune the filter module gain to a few %. Seems good enough for now, but we should really try to understand what it is and why it is the way that it is. In the above plot, the ORANGE trace is the old cancellation and the GREEN one is the new one. The filter TF is attached below - its not special, we made it by presing buttons in FOTON until the TF matched the measured TF of ALSY/LSC-MC_CTRL_FF_OUT.

Attachment 1: ALSfoo_150216.png
ALSfoo_150216.png
Attachment 2: 15.png
15.png
  11040   Mon Feb 16 21:52:51 2015 ranaHowToTreasurebig Dataviewer windows

Following this entry, I have made the same change in the controls account on rossa:

In the ~/.grace/gracerc file (create one if it doesn't exist), put in a line which reads:

PAGE LAYOUT FREE

Now we can scale our dataviewer live and playback plots by stretching the window with our mouse. The attached screenshot shows how I filled up one of the vertical monitors with a DV window for arm locking.

Attachment 1: bigDV.png
bigDV.png
  11039   Mon Feb 16 15:08:26 2015 JenneUpdateLSCALS fool measured decoupling TF

Dang it, yes. You're right.  I should have caught that. 

Since Dcpl and Srefl are both zero during the measurement (since it was an ALS lock), this is actually

\frac{{\color{DarkRed} A_{\rm refl}} {\color{DarkGreen} P_{\rm als} S_{\rm als}}}{1 - {\color{DarkGreen} A_{\rm als} G_{\rm als} S_{\rm als} P_{\rm als}}}

So, I need to remove the effect of the ALS closed loop, to get the actual quantity I was looking for.

  11038   Mon Feb 16 03:10:42 2015 KojiUpdateLSCALS fool measured decoupling TF

Wonkey shape: Looks like a loop supression. Your http://nodus.ligo.caltech.edu:8080/40m/11016 also suggests it too, doesn't it?

  11037   Mon Feb 16 02:49:57 2015 JenneUpdateLSCALS fool measured decoupling TF

I have measured very, very carefully the transfer function from pushing on MC2 to the Yarm ALS beatnote.  In the newest loop diagram in http://nodus.ligo.caltech.edu:8080/40m/11030, this is pushing at point 10 and sensing at point 4. 

Since it's a bunch of different transfer functions (to get the high coherence that we need for good cancellation to be possible), I attach my Matlab figure that includes only the useful data points.  I put a coherence cutoff of 0.99, so that (assuming the fit were perfect, which it won't be), we would be able to get a maximum cancellation of a factor of 100. 

This plot also includes the vectfit to the data, which you can see is pretty good, although I need to separately plot the residuals (since the magnitude data is so small, the residuals for the mag don't show up in the auto plot that vectfit gives). 

If you recall from http://nodus.ligo.caltech.edu:8080/40m/11020, we are expecting this transfer function to consist of the suspension actuator (pendulum with complex pole pair around 1Hz), the ALS plant (single pole at 80kHz) and the ALS sensor shape (the phase tracker is an integrator, with a boost consisting of a zero at 666Hz and a pole at 55Hz).  That expected transfer function does not multiply up to give me this wonky shape.  Brain power is needed here.

Just in case you were wondering if this depends on the actuator used (ETM vs MC2), or IFO configuration (single arm vs. PRFPMI), it doesn't.  The only discrepancy between these transfer functions is the expected sign flip between the MC2 and ETMY actuators.  So, they're all pretty consistent. 

Before locking the PRFPMI, I copied the boost filter (666:55) from the YARM ALS over to Xarm ALS, so now both arms have the same boost.

YARM_actTF_compareActuators.pdf


Things to do for ALSfool:

  • Put fitted TF into the MC_CTRL_FF filter bank, and try to measure the expected cancellation, a la http://nodus.ligo.caltech.edu:8080/40m/11009
  • Quick test with single arm, ALS locked using full loop (high gain, all boosts), since the previous versions were with ALS very loosely locked.
    • Does this measured transfer function actually give us good cancellation? 
  • Think.  Why should the transfer function look like this??
  • Try it on the full PRFPMI
Attachment 1: ALSfool_measuredActuatorTF_YarmOnly_15Feb2015.png
ALSfool_measuredActuatorTF_YarmOnly_15Feb2015.png
Attachment 2: YARM_actTF_compareActuators.pdf
YARM_actTF_compareActuators.pdf
  11036   Mon Feb 16 01:45:12 2015 KojiSummaryLSCmodulation depth analysis

Based on the measured modulation profiles, the depth of each modulation was estimated.
Least square sum minimization of the relative error was used for the cost function.
-8th, -12th~-14th, n=>7th are not included in the estimation for the nominal case.
-7th~-9th, -11th~-15th, n=>7th are not included in the estimation for the 3f reduced case.

Nominal modulation

m_f1 = 0.194
m_f2 = 0.234
theta_f1f2 = 41.35deg
m_IMC = 0.00153

3f reduced modulation

m_f1 = 0.191
m_f2 = 0.0579
theta_f1f2 = 180deg
m_IMC = 0.00149

(Sorry! There is no error bars. The data have too few statistics...)

Attachment 1: modulation_nominal.pdf
modulation_nominal.pdf
Attachment 2: modulation_3f_reduced.pdf
modulation_3f_reduced.pdf
  11035   Mon Feb 16 00:08:44 2015 KojiSummaryLSC[ELOG LIST] 3f modulation cancellation

This KTP crystal has the maximum allowed RF power of 10W (=32Vpk) and V_pi = 230V. This corresponds to the maximum allowed
modulation depth of 32*Pi/230 = 0.44. So we probably can achieve gamma_1 of ~0.4 and gamma_2 of ~0.13. That's not x3 but x2,
so sounds not too bad.

Then Kiwamu's triple resonant circuit LIGO-G1000297-v1 actually shows the modulation up to ~0.7. Therefore it is purely an issue
how to deliver sufficient modulation power. (In fact his measurement shows some nonlinearity above the modulation depth of ~0.4
so we should keep the maximum power consumption of 10W at the crystal)

This means that we need to review our RF system (again!)

- Review infamous crazy attn/amp combinations in the frequency generation box.
- Use Teledyne Cougar ampilfier (A2CP2596) right before the triple resonant box. This should be installed closely to the triple resonant box in order to
minimize the effects of the reflection due to imperferct impedance matching.
- Review and refine the triple resonant circuit - it's not built on a PCB but on a universal board. I think that we don't need triple
resonance, but double is OK as the 29.5MHz signal is small.

We want +28V supply at 1X1 for the Teledyne amp and the AOM driver. Do we have any unused Sorensen?

  11034   Sun Feb 15 20:55:48 2015 ranaSummaryLSC[ELOG LIST] 3f modulation cancellation

I wonder if DRMI can be locked on 3f using this lower 55 MHz modulation depth. It seems that PRMI should be unaffected, but that the 3*f2 signals for SRCL will be too puny. Is it really possible to scale up the overall modulation depths by 3x to compensate for this?

  11033   Sun Feb 15 16:20:44 2015 KojiSummaryLSC[ELOG LIST] 3f modulation cancellation

Summary of the ELOGS

3f modulation cancellation theory http://nodus.ligo.caltech.edu:8080/40m/11005

3f modulation cancellation adjustment setup http://nodus.ligo.caltech.edu:8080/40m/11029

Experiment http://nodus.ligo.caltech.edu:8080/40m/11031

Receipe for the 3f modulation cancellation http://nodus.ligo.caltech.edu:8080/40m/11032

Modulation depth analysis http://nodus.ligo.caltech.edu:8080/40m/11036

  11032   Sat Feb 14 22:14:02 2015 KojiSummaryLSC[HOW TO] 3f modulation cancellation

When I finished my measurements, the modulation setup was reverted to the conventional one.
If someone wants to use the 3f cancellation setting, it can be done along with this HOW-TO.


The 3f cancellation can be realized by adding a carefully adjusted delay line and attenuation for the 55MHz modulation
on the frequency generation box at the 1X2 rack.  Here is the procedure:

1) Turn off the frequency generation box

There is a toggle switch at the rear of the unit. It's better to turn it off before any cable action.
The outputs of the frequency generation box are high in general. We don't want to operate
the amplifiers without proper impedance matching in any occasion.

2) Remove the small SMA cable between 55MHz out and 55MHz in (Left arrow in the attachment 1).

According to the photo by Alberto (svn: /docs/upgrade08/RFsystem/frequencyGenerationBox/photos/DSC_2410.JPG),
this 55MHz out is the output of the frequency multiplier. The 55MHz in is the input for the amplifier stages.
Therefore, the cable length between these two connectors changes the relative phase between the modulations at 11MHz and 55MHz.

3) Add a delay line box with cables (Attachment 2).

Connect the cables from the delay line box to the 55MHz in/out connectors. I used 1.5m BNC cables.
The delay line box was set to have 28ns delay.

4) Set the attenuation of the 55MHz EOM drive (Right arrow in the attachment 1) to be 10dB.

Rotate the attenuation for 55MHz EOM from 0dB nominal to 10dB.

5) Turn on the frequency modulation box


For reference, the 3rd attachment shows the characteristics of the delay line cable/box combo when the 3f modualtion reduction
was realized. It had 1.37dB attenuation and +124deg phase shift. This phase change corresponds to the time delay of 48ns.
Note that the response of a short cable used for the measurement has been calibrated out using the CAL function of the network analyzer.

Attachment 1: freq_gen_box.JPG
freq_gen_box.JPG
Attachment 2: delay_line.JPG
delay_line.JPG
Attachment 3: cable_spec.pdf
cable_spec.pdf
  11031   Sat Feb 14 20:37:51 2015 KojiSummaryLSC3f modulation cancellation

Experimental results

- PD response [Attachment 1]

The AUX laser temperature was swept along with the note by Annalisa [http://nodus.ligo.caltech.edu:8080/40m/8369]
It is easier to observe the beat note by closing the PSL shutter as the MC locking yields more fluctuation of the PSL
laser freuqency at low frequency. Once I got the beat note and maximized it, I immediately noticed that the PD response
is not flat. For the next trial, we should use Newfocus 1611. For the measurement today, I decided to characterize the
response by sweeping the beat frequency and use the MAXHOLD function of the spectrum analyzer.

The measured and modelled response of the PD are shown in the attachment 1. It has non-intuitive shape.
Therefore the response is first modelled by two complex pole pair at 127.5MHz with Q of 1, and then the residual was
empirically fitted with 29th polynomial of f.

- Modulation profile of the nominal setting [Attachment 2]

Now the spectrum of the PD output was measured. This is a stiched data of the spectrum between 1~101MHz and 99~199MHz
that was almost simultaneously measured (i.e. Display 1 and Display 2). The IF bandwidth was 1kHz. The PD response correction
described above was applied.

It obviously had the peaks associated with our main modulations. In addition, there are more peaks seen.
The attachment 2 breaks down what is causing the peaks.

  • Carrier: The PLL LO frequency is 95MHz. Therefore the carrier is locked at 95MHz.
  • Modulation sidebands (11/55MHz series):
    Series of sidebands are seen at the both side of the carrier. Their frequency is 95MHz +/- n * fmod  (fmod = 11.066128MHz).
    Note that the sidebands for n>10 were above 200MHz, and n<-9 (indicated in gray) were folded at 0Hz.
    With this measurement BW, the following sidebands were buried in the noise floor.
    n = -8, -12, -13, and -14, n<= -16, and n>=+7
  • Modulation sidebands for IMC and PMC (29.5MHz and 35.5MHz):
    First order sidebands for the IMC and PMC modulations of sidebands are seen at the both side of the carrier.
    Their frequency is 95MHz +/- 29.5MHz or 33.5MHz. The PMC modulation sidebands are supposed to be blocked
    by the PMC. However, due to finite finesse of the PMC, small fraction of the PMC sidebands are transmitted.
    In deed, it is comparable to the modulation depth of the IMC one.
  • RF AM or RF EMI for the main modulation and the IMC modulationand:
    If there is residual RF AM in the PSL beam associated with the IMC and main modulations, it appears as the
    peaks at the modulation frequency and its harmonics. Also EM radiation couples into this measument RF system
    also appears at these frequencies. They are seen at n * fmod  (n=1,2,4,5) and 29.5MHz.
  • Reflection/distortion or leakage from mixer IF to RF:
    The IF port of the mixer naturally has 190MHz signal when the PLL is locked. If the isolation from the IF port to the RF port
    is not enough, this signal can appear in the RF monitor signal via an imperfection of the coupler or a reflection from the PD.
    Also, if the reflecrtion/distortion exist between the PD and the mixer RF input, it also cause the signal around 190MHz.
    It is seen at 190MHz +/- n* fmod. In the plot, the peak at n=0, -1 are visible. In fact these peak were secondarily dominant
    in the spectrum when there was no 6dB attenuation in the PD line. WIth the attenuator, they are well damped and don't disturb
    the main measurment.

From the measured peak height, we are able to estimate the modulation depths for 11MHz, 55MHz, IMC modulations, as well as
the relative phase of the 11MHz and 55MHz modulation. (It is not yet done).

- 3f modulation reduction [Attachment 3]

Now, the redcution of the 3f modulation was tried. The measured modulation levels for the 11MHz and 55MHz were almost the same.
The calculation predicts that the modulation for the 55MHz needs to be 1/3 of the 11MHz one. Therefore the attenuation of 9dB and 10dB
of the modulation attenuation knob at the frequency generation box were tried.

To give the variable delay time in the 55MHz line, EG&G ORTEC delay line unit was used. This allows us to change the delay time from
0ns to 63.5ns with the resolution of 0.5ns. The frequency of 55MHz yields a phase sensitivity of ~20deg/ns (360deg/18ns).
Therefore we can adjust the phase with the precision of 10deg over 1275deg.

The 3rd-order peak at 61.8MHz was observed with measurement span of 1kHz with very narrow BW like 30Hz(? not so sure). The delay
time was swept while measuring the peak height each time. For both the atteuation, the peak height clearly showed the repeatitive dependence
with the period of 18ns, and the 10dB case gave the better result. The difference between the best (1.24e-7 Vpk) and the worst (2.63e-6 Vpk)
was more than a factor of 20.
The 3rd-order peak in the above broadband spectrum measurement was 6.38e-6 Vpk. Considering the attenuation
of the 55MHz modulation by 10dB, we were at the exact unluck phase difference.
The improvement expected from the 3f reduction (in the 33MHz signal)
will be about 50, assuming there is no other coupling mechanism from CARM to REFL33.

I decided to declare the best setting is "10dB attenuation & 28ns delay".

- Resulting modulation profile [Attachment 4]

As a confirmation, the modulation profie was measured as done before the adjustment.
It is clear that the 3rd-order modulation was buried in the floor noise. 10dB attenuation of the 55MHz modulation yields corresponding reduction of the sidebands.
This will impact the signal quality for the 55MHz series error signals, particularly 165MHz ones. We should consider to install the Teledyne Cougar amplifier
next to the EOM so that we can increase the over all modulation depth.

Attachment 1: beat_pd_response.pdf
beat_pd_response.pdf
Attachment 2: beat_nominal.pdf
beat_nominal.pdf
Attachment 3: 3f_reduction.pdf
3f_reduction.pdf
Attachment 4: beat_3f_reduced.pdf
beat_3f_reduced.pdf
  11030   Sat Feb 14 20:20:24 2015 JenneUpdateLSCALS fool cartoon

The ALS fool scheme is now diagrammed up in OmniGraffle, including its new official icon.  The mathematica notebook has not yet been updated.

EDIT, JCD, 17Feb2015:  Updated cartoon and calculation: http://131.215.115.52:8080/40m/11043

 

Attachment 1: ALSfool_LoopDiagram.png
ALSfool_LoopDiagram.png
Attachment 2: ALSfool_LoopDiagram.graffle.zip
  11029   Sat Feb 14 19:54:04 2015 KojiSummaryLSC3f modulation cancellation

Optical Setup

[Attachment 1]

Right before the PSL beam goes into the vacuum chamber, it goes through an AR-wedged plate.
This AR plate produces two beams. One of them is for the IO beam angle/position monitor.
And the other was usually dumped. I decided to use this beam.

A G&H mirror reflects the beam towards the edge of the table.
A 45deg HR mirror brings this beam to the beat set up at the south side of the table.
This beam is S-polarlized as it directly comes from the EOM.

[Attachment 2]

The beam from the PSL goes through a HWP and some matching lenses before the combining beam splitter (50% 45deg P).
The AUX laser beam is attenuated by a HWP and a PBS. The transmitted beam from the PBS is supposed
to have P-polarization. The beam alignment is usually done at the PSL beam side.

The combined beam is steered by a HR mirror and introduced to Thorlabs PDA10CF. As the PD has small diameter
of 0.5mm, the beam needed to be focused by a strong lens.

After careful adjustment of the beam mode matching, polarization, and alignment, the beatnote was ~1Vpp for 2.5Vdc.
In the end, I reduced the AUX laser power such that the beat amplitude went down to ~0.18Vpp (-11dBm at the PD,
-18dBm at the mixer, -27dBm at the spectrum analyzer) in order to minimize nonlinearity of the RF system and
in order that the spectrum analyzer didn't need input attenuation.

Electrical Setup

[Attachment 3]

The PD signal is mixed with a local oscillator signal at 95MHz, and then used to lock the PLL loop.
The PLL loop allows us to observe the peaks with more integration time, and thus with a better signal-to-noise ratio.

The signal from the PD output goes through a DC block, then 6dB attenuator. This attenuator is added to damp reflection
and distortion between the PD and the mixer. When the PLL is locked, the dominant signal is the one at 95MHz. Without this attenuator,
this strong 95MHz signal cause harmonic distortions like 190MHz. As a result, it causes series of spurious peaks at 190MHz +/- n* 11MHz.

10dB coupler is used to peep the PD signal without much disturbing the main line. Considering we have 6dB attanuator,
we can use this coupler output for the PLL and can use the main line for the RF monitor, next time.

The mixer takes the PD signal and the LO signal from Marconi. Marconi is set to have +7dBm output at 95MHz.
FOr the image rejection, SLP1.9 was used. The minicirsuit filters have high-Z at the stop band, we need a 50Ohm temrinator
between the mixer and the LPF.

The error signal from the LPF is fed to SR560 (G=+500, 1Hz 1st-order LPF). I still don't understand why I had to use a LPF
for the locking.
As the NPRO PZT is a frequency actuator, and the PLL is sensitive to the phase, we are supposed to use
a flat response for PLL locking. But it didn't work. Once we check the open loop TF of the system, it will become obvious (but I didn't).

The actuation signal is fed to the fast PZT input of the AUX NPRO laser.
 

Attachment 1: beat_setup1.JPG
beat_setup1.JPG
Attachment 2: beat_setup2.JPG
beat_setup2.JPG
Attachment 3: electrical_setup.pdf
electrical_setup.pdf
  11028   Sat Feb 14 00:48:13 2015 KojiUpdateLSC3f modulation cancellation

[SUCCESS] The 3f sideband cancellation seemed worked nicely.

- Beat effeciency improved: ~30% contrast (no need for amplification)

- PLL locked

- 3f modulation sideband was seen

- The attenuation of the 55MHz modulation and the delay time between the modulation source was adjusted to
have maximum reduction of the 3f sidebands as much as allowed in the setup. This adjustment has been done
at the frequency generation box at 1X2 rack.

- The measurement and receipe for the sideband cancellation come later.


- This means that I jiggled the modulation setup at 1X2 rack. Now the modulation setup was reverted to the original,
but just be careful to any change of the sensing behavior.

- The RF analyzer was returned to the control room.

- The HEPA speed was reduced from 100% (during the action on the table) to 40%.

  11027   Sat Feb 14 00:42:02 2015 KojiUpdateGeneralRF amplifier for ALS

The RF analyzer was returned to the control room. There are two beat notes from X/Y confirmed.

I locked the arms and aligned them with ASS.

When the end greens are locked at TEM00, X/Y beat amplitudes were ~33dBm and ~17dBm. respectively.
I don't judge if they are OK or not, as I don't recall the nominal values.

  11026   Fri Feb 13 19:41:13 2015 ericqUpdateGeneralRF amplifier for ALS

As Koji found one of the spare channels of the ALS/FOL RF amplifier box nonfunctional yesterday, I pulled it out to fix it. I found that one of the sma cables did not conduct.

It was replaced with a new cable from Koji. Also, I rearranged the ports to be consistent across the box, and re-labeled with the gains I observed. 

It has been reinstalled, and the Y frequency counter that is using one of the channels shows a steady beat freq.

I cannot test the amplitude of the green X beat at this time, as Koji is on the PSL table with the PSL shutter closed, and is using the control room spectrum analyzer. However, the dataviewer trace for the fine_phase_out_Hz looks like free swinging cavity motion, so its probably ok.

  11025   Fri Feb 13 18:56:44 2015 ranaUpdateElectronicsSecond QPD Whitening Switch enabled

Depends on the plots of the whitening I guess; if its low freq sat, then we lower the light level with ND filters. If its happening above 10 Hz, then we switch off the whitening.

Quote:

Went to zero CARM offset on ALS; transmission QPDs are still saturating :(

Maybe we need to switch off all whitening.

 

  11024   Fri Feb 13 17:07:51 2015 JenneUpdateElectronicsSecond QPD Whitening Switch enabled

I first updated the DCC branches for the Xend and Yend to reflect the as-built situation from December 2014, and then I updated the drawings after Q's modifications today.

  11023   Fri Feb 13 14:59:13 2015 ericqUpdateElectronicsSecond QPD Whitening Switch enabled

Went to zero CARM offset on ALS; transmission QPDs are still saturating :(

Maybe we need to switch off all whitening.

  11022   Fri Feb 13 14:27:30 2015 ericqUpdateElectronicsSecond QPD Whitening Switch enabled

I have re-enabled the second whitening stage switching on each quadrant of each end's QPD whitening board, to try and avoid saturations at full power. Looking at the spectra while single arm locked, I confirmed that the FM2 whitening switch works as expected. FM1 should be left on, as it is still hard-wired to whiten. 

The oscillations in the Y QPD still exist. Jenne is updating the schematics on the DCC.

  11021   Fri Feb 13 03:44:56 2015 JenneUpdateLSCHeld using ALS for a while at "0" CARM offset with PRMI

[Jenne, Rana]

We wanted to jump right in and see if we were ready to try the new "ALS fool" loop decoupling scheme, so we spent some time with CARM and DARM at "0" offset, held on ALS, with PRMI locked on REFL33I&Q (no offsets).  Spoiler alert:  we weren't ready for the jump.

The REFL11 and AS55 PDs had 0dB analog whitening, which means that we weren't well-matching our noise levels between the PD noise and the ADC noise.  The photodiodes have something of the order nanovolt level noise, while the ADC has something of the order microvolt level noise.  So, we expect to need an analog gain of 1000 somewhere, to make these match up.  Anyhow, we have set both REFL11 and AS55 to 18dB gain. 

On a related note, it seems not so great for the POX and POY ADC channels to be constantly saturated when we have some recycling gain, so we turned their analog gains down from 45dB to 0dB.  After we finished with full IFO locking, they were returned to their nominal 45dB levels. 

We also checked the REFL33 demod phase at a variety of CARM offsets, and we see that perhaps it changes by one or two degrees for optimal rotation, but it's not changing drastically.  So, we can set the REFL33 demod phase at large CARM offset, and trust it at small CARM offset.

We then had a look at the transmon QPD inputs (before the dewhitening) for each quadrant.  They are super-duper saturating, which is not so excellent. 
QPDsaturation_12Feb2015.pdf

We think that we want to undo the permanently-on whitening situation.  We want to make the second stage of whitening back to being switchable.  This means taking out the little u-shaped wires that are pulling the logic input of the switches to ground.  We think that we should be okay with one always on, and one switchable.  After the modification, we must check to make sure that the switching behaves as expected.  Also, I need to figure out what the current situation is for the end QPDs, and make sure that the DCC document tree matches reality.  In particular, the Yend DCC leaf doesn't include the gain changes, and the Xend leaf which does show those changes has the wrong value for the gain resistor.

After this, we started re-looking at the single arm cancellation, as Rana elogged about separately.

ALSfool_12Feb2015.pdf

Attachment 1: QPDsaturation_12Feb2015.pdf
QPDsaturation_12Feb2015.pdf
Attachment 2: ALSfool_12Feb2015.pdf
ALSfool_12Feb2015.pdf
  11020   Fri Feb 13 03:28:34 2015 ranaUpdateLSCNew Locking Paradigm - Loop-gebra

Not so fast!

In the drawing, the FF path should actually be summed in after the Phase Tracker (i.e. after S_ALS). This means that the slow response of the phase tracker needs to be taken into account in the FF cancellation filter. i.e. D = -A_REFL * P_ALS * S_ALS. Since the Phase Tracker is a 1/f loop with a 1 kHz UGF, at 100 Hz, we can only get a cancellation factor of ~10.

So, tonight we added a 666:55 boost filter into the phase tracker filter bank. I think this might even make the ALS locking loops less laggy. The boost is made to give us better tracking below ~200 Hz where we want better phase performance in the ALS and more cancellation of the ALS-Fool. If it seems to work out well we can keep it. If it makes ALS more buggy, we can just shut it off.

Its time to take this loop cartoon into OmniGraffle.

  11019   Thu Feb 12 23:47:45 2015 KojiUpdateLSC3f modulation cancellation

- I built another beat setup on the PSL table at the South East side of the table.
- The main beam is not touched, no RF signal is touched, but recognize that I was present at the PSL table.
- The beat note is found. The 3rd order sideband was not seen so far.
- A PLL will be built tomorrow. The amplifier box Manasa made will be inspected tomorrow.

- One of the two beams from the picked-off beam from the main beam line was introduced to the beat setup.
(The other beam is used of for the beam pointing monitors)
There is another laser at that corner and the output from this beam is introduced into the beat setup.
The combined beam is introduced to PDA10CF (~150MHz BW).

- The matching of the beam there is poor. But without much effort I found the beat note.
  The PSL laser had 31.33 deg Xtal temp. When the beat was found, the aux laser had the Xtal temp of 40.88.

- I could observe the sidebands easily, with a narrower BW of the RF analizer I could see the sidebands up to the 2nd order.
  The 3rd order was not seen at all.

- The beat note had the amplitude of about -30dBm. One possibility is to amplify the signal. I wanted to use a spare channel
of the ALS/FOLL amplifier box. But it gave me rather attenuation than any amplification.
I'll look at the box tomorrow.

- Also the matching of two beams are not great. The PD also has clipping I guess. These will also be improved tomorrow

- Then the beat note will be locked at a certain frequency using PLL so that we can reduce the measurement BW more.

 

  11018   Thu Feb 12 23:40:12 2015 ranaUpdateIOOBounce / Roll stopband filters added to MC ASC filter banks

The filters were already in the damping loops but missing the MC WFS path. I checked that these accurately cover the peaks at 16.5 Hz and 23.90 and 24.06 Hz.

Attachment 1: 59.png
59.png
  11017   Thu Feb 12 22:28:16 2015 JenneUpdateLSCNew Locking Paradigm - LSC model changes, screens modified

I have modified the LSC trigger matrix screen, as well as the LSC overview screen, to reflect the modifications to the model from yesterday. 

Also, I decided that we probably won't ever want to trigger the zero crossing on the Q phase signals of REFL.  Instead, we may want to try it out with the single arms, so the zero crossing selection matrix is now REFL11I, REFL55I, POX11I, POY11I, in that order. 

 

  11016   Thu Feb 12 19:18:49 2015 JenneUpdateLSCNew Locking Paradigm - Loop-gebra

EDIT, JCD, 17Feb2015:  Updated loop diagram and calculation: http://131.215.115.52:8080/40m/11043


Okay, Koji and I talked (after he talked to Rana), and I re-looked at the original cartoon from when Rana and I were thinking about this the other day.

The original idea was to be able to actuate on the MC frequency (using REFL as the sensor), without affecting the ALS loop.  Since actuating on the MC will move the PSL frequency around, we need to tell the ALS error signal how much the PSL moved in order to subtract away this effect. (In reality, it doesn't matter if we're actuating on the MC or the ETMs, but it's easier for me to think about this way around).  This means that we want to be able to actuate from point 10 in the diagram, and not feel anything at point 4 in the diagram (diagram from http://131.215.115.52:8080/40m/11011)

This is the same as saying that we wanted the green trace in http://131.215.115.52:8080/40m/11009 to be zero.

So.  What is the total TF from 10 to 4? 

{\rm TF}_{\rm (10 \ to \ 4)} = \frac{D_{\rm cpl} + {\color{DarkRed} A_{\rm refl}} {\color{DarkGreen} P_{\rm als}}}{1-{\color{DarkRed} A_{\rm refl} G_{\rm refl} S_{\rm refl} P_{\rm refl}} - {\color{DarkGreen} A_{\rm als} G_{\rm als} S_{\rm als}} ({\color{DarkGreen} P_{\rm als}} + D_{\rm cpl} {\color{DarkRed} G_{\rm refl} P_{\rm refl} S_{\rm refl}})}

So, to set this equal to zero (ALS is immune to any REFL loop actuation), we need D_{\rm cpl} = - {\color{DarkRed} A_{\rm refl}} {\color{DarkGreen} P_{\rm als}}.

Next up, we want to see what this means for the closed loop gain of the whole system.  For simplicity, let's let H_* = A_* G_* S_* P_*, where * can be either REFL or ALS. 

Recall that the closed loop gain of the system (from point 1 to point 2)  is

{\rm TF}_{\rm (1 \ to \ 2)} = \frac{1}{1-{\color{DarkRed} A_{\rm refl} G_{\rm refl} S_{\rm refl} P_{\rm refl}} - {\color{DarkGreen} A_{\rm als} G_{\rm als} S_{\rm als}} ({\color{DarkGreen} P_{\rm als}} + D_{\rm cpl} {\color{DarkRed} G_{\rm refl} P_{\rm refl} S_{\rm refl}})} , so if we let  D_{\rm cpl} = - {\color{DarkRed} A_{\rm refl}} {\color{DarkGreen} P_{\rm als}} and simplify, we get

{\rm TF}_{\rm (1 \ to \ 2)} = \frac{1}{1-{\color{DarkRed} H_{\rm refl}} - {\color{DarkGreen} H_{\rm als}} + {\color{DarkRed} H_{\rm refl}}{\color{DarkGreen} H_{\rm als}}}

This seems a little scary, in that maybe we have to be careful about keeping the system stable.  Hmmmm.  Note to self:  more brain energy here.


Also, this means that I cannot explain why the filter wasn't working last night, with the guess of a complex pole pair at 1Hz for the MC actuator.  The  ALS plant has a cavity pole at ~80kHz, so for our purposes is totally flat.  The only other thing that comes to mind is the delays that exist because the ALS signals have to hop from computer to computer.  But, as Rana points out, this isn't really all that much phase delay below 100Hz where we want the cancellation to be awesome. 

I propose that we just measure and vectfit the transfer function that we need, since that seems less time consuming than iteratively tweaking and checking. 

Also, I just now looked at the wiki, and the MC2 suspension resonance for pos is at 0.97Hz, although I don't suspect that that will have changed anything significantly above a few Hz.  Maybe it makes the cancellation right near 1Hz a little worse, but not well above the resonance.

 

  11015   Thu Feb 12 15:21:37 2015 ericqUpdateComputer Scripts / Programsnetgpib updates

I've fixed the gpib scripts for the SR785 and AG4395A to output data in the same format as expected by older scripts when called by them. In addition, there are now some easier modes of operation through the measurement scripts SRmeasure and AGmeasure. These are on the $PATH for the main control room machines, and live in scripts/general/netgpib


Case 1: I manually set up a measurement on the analyzer, and just want to download / plot the data. 

Make sure you have a yellow prologix box plugged in, and can ping the address it is labeled with. (i.e. 'vanna'). Then, in the directory you want to save the data, run:

SRmeasure -i vanna -f mydata --getdata --plot

This saves mydata_(datetime).txt and mydata_(datetime).pdf in the current directory. 

In all cases, AGmeasure has the identical syntax. If the GPIB address is something other than 10, specifiy it with -a, but this is rarely the case.


Case 2: I want to remotely specify a measurement

Rather than a series of command line arguments, which may get lost to the mists of time, I've set the scripts up to use parameter files that serve as arguments to the scripts. 

Get the templates for spectrum and TF measurements in your current directory by running

SRmeasure --template

Set the parameters with your text editor of choice, such as frequency span, filename output, whether to create a plot or not, then run the measurement:

SRmeasure SR785template.yml


Case 3: I want to compare my data with previous measurements

In the template parameter files, there is an option 'plotRefs', that will automatically plot the data from files whose filenames start with the same string as the current measurement.

If, in the "#" commented out header of the data file, there is a line that contains "memo:" or "timestamp:", it will include the text that follows in the plot legend. 


There are also methods to remotely trigger an already configured measurement, or remotely reset an unresponsive instrument. Options can be perused by looking at the help in SRmeasure -h

I've tested, debugged, and used them for a bit, but wrinkles may remain. They've been svn40m committed, and I also set up a separate git repository for them at github.com/e-q/netgpibdata

  11014   Thu Feb 12 12:23:21 2015 manasaUpdateGeneralWaking up the PDFR measurement system

[EricG, Manasa]

We woke up the PDFR measurement setup that has been sleeping since summer. We ran a check for the laser module and the multiplexer module. We tried setting things up for measuring frequency response of AS55.
We could not repeat Nichin's measurements because the gpib scripts are outdated and need to be revised. 

PDFR diode laser was shutdown after this job.

  11013   Thu Feb 12 12:16:04 2015 manasaUpdateGeneralFiber shielding

[Steve, Manasa]

The fibers around the PSL table were shielded to avoid any tampering.

  11012   Thu Feb 12 11:59:58 2015 KojiUpdateLSCNew Locking Paradigm - Loop-gebra

The goals are:

- When the REFL path is dead (e.g. S_REFL = 0), the system goes back to the ordinary ALS loop. => True (Good)

- When the REFL path is working, the system becomes insensityve to the ALS loop
(i.e. The ALS loop is inactivated without turning off the loop.) => True when (...) = 0

Are they correct?

 

Then I just repeat the same question as yesterday:

S is a constant, and Ps are cavity poles. So,  approximately to say, (...) = 0 is realized by making D = 1/G_REFL.
In fact, if we tap the D-path before the G_REFL, we remove this G_REFL from (...). (=simpler)
But then, this means that the method is rather cancellation between the error signals than
cancellation between the actuation. Is this intuitively reasonable? Or my goal above is wrong?

  11011   Thu Feb 12 11:14:29 2015 JenneUpdateLSCNew Locking Paradigm - Loop-gebra

I have calculated the response of this new 2.5 loop system.

The first attachment is my block diagram of the system.  In the bottom left corner are the one-hop responses from each green-colored point to the next.  I use the same matrix formalism that we use for Optickle, which Rana described in the loop-ology context in http://nodus.ligo.caltech.edu:8080/40m/10899

In the bottom right corner is the closed loop response of the whole system.

 

Also attached is a zipped version of the mathematica notebook used to do the calculation.

EDIT, JCD, 17Feb2015:  Updated loop diagram and calculation:  http://131.215.115.52:8080/40m/11043

Attachment 1: ALS_REFL_comboLockingCartoon_11Feb2015.PDF
ALS_REFL_comboLockingCartoon_11Feb2015.PDF
Attachment 2: ALS_REFL_comboLocking_11Feb2015.zip
  11010   Thu Feb 12 03:43:54 2015 ericqUpdateLSC3F PRMI at zero ALS CARM

I have been able to recover the ability to sit at zero CARM offset while the PRMI is locked on RELF33 and CARM/DARM are on ALS, effectively indefinitely. However, I feel like the transmon QPDs are not behaving ideally, because the reported arm powers freqently go negative as the interferometer is "buzzing" through resonance, so I'm not sure how useful they'll be as normalizing signals for REFL11. I tried tweaking the DARM offset to help the buildup, since ALS is only roughly centered on zero for both CARM and DARM, but didn't have much luck.

Example:

Turning off the whitening on the QPD segments seems to make everything saturate, so some thinking with daytime brain is in order.


How I got there:

It turns out triggering is more important than the phase margin story I had been telling myself. Also, I lost a lot of time to needing demod angle change in REFL33. Maybe I somehow caused this when I was all up on the LSC rack today?

We have previously put TRX and TRY triggering elements into the PRCL and MICH rows, to guard against temporary POP22 dips, because if arm powers are greater than 1, power recylcing is happening, so we should keep the loops engaged. However, since TRX and TRY are going negative when we buzz back and forth through the resonsnace, the trigger row sums to a negative value, and the PRMI loops give up. 

Instead, we can used the fortuitously unwhitened POPDC, which can serve the same function, and does not have the tendancy to go negative. Once I enabled this, I was able to just sit there as the IFO angrily buzzed at me. 

Here are my PRMI settings

REFL33 - Rotation 140.2 Degrees, -89.794 measured diff

PRCL = 1 x REFL33 I; G = -0.03; Acquire FMs 4,5; Trigger FMs 2, 9; Limit: 15k ; Acutate 1 x PRM

MICH = 1 x REFL33 Q, G= 3.0, Acquire FMs 4,5,8; Trigger FM 2, 3; Limit: 30k; Actuate -0.2625 x PRM + 0.5 x BS

Triggers = 1 x POP22 I + 0.1 * POPDC, 50 up 5 down


Just for kicks, here's a video of the buzzing as experienced in the control room

Attachment 1: Feb12_negativeTR.png
Feb12_negativeTR.png
  11009   Thu Feb 12 01:43:09 2015 ranaUpdateLSCNew Locking Paradigm - LSC model changes

With the Y Arm locked, we checked that we indeed can get loop decoupling using this technique.

The guess filter that we plugged in is a complex pole pair at 1 Hz. We guessed that the DC gain should be ~4.5 nm count. We then converted this number into Hz and then into deg(?) using some of Jenne's secret numbers. Then after measuring, we had to increase this number by 14.3 dB to an overall filter module gain of +9.3.

The RED trace is the usual 'open loop gain' measurement we make, but this time just on the LSC-MC path (which is the POY11_I -> ETMY path).

The BLUE trace is the TF between the ALS-Y phase tracker output and the FF cancellation signal. We want this to be equal ideally.

The GREEN trace is after the summing point of the ALS and the FF. So this would go to zero when the cancellation is perfect.

So, not bad for a first try. Looks like its good at DC and worse near the red loop UGF. It doesn't change much if I turn off the ALS loop (which I was running with ~10-15x lower than nominal gain just to keep it out of the picture). We need Jenne to think about the loop algebra a little more and give us our next filter shape iteration and then we should be good.

Attachment 1: TF.gif
TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif TF.gif
  11008   Thu Feb 12 01:00:18 2015 ranaUpdateLSCRFPD spectra

The nonlinearity in the LSC detection chain (cf T050268) comes from the photodetector and not the demod board. The demod board has low pass or band pass filters which Suresh installed a long time ago (we should check out what's in REFL33 demod board). 

Inside the photodetector the nonlinearity comes about because of photodiode bias modulation (aka the Grote effect) and slew rate limited distortion in the MAX4107 preamp.

  11007   Wed Feb 11 22:13:44 2015 JenneUpdateLSCNew Locking Paradigm - LSC model changes

In order to try out the new locking scheme tonight, I have modified the LSC model.  Screens have not yet been made.

It's a bit of a special case, so you must use the appropriate filter banks:

CARM filter bank should be used for ALS lock.  MC filter bank should be used for the REFL1f signal. 

The output of the MC filter bank is fed to a new filter bank (C1:LSC-MC_CTRL_FF).  The output of this new filter bank is summed with the error point of the CARM filter bank (after the CARM triggered switch).

The MC triggering situation is now a little more sophisticated than it was.  The old trigger is still there (which will be used for something like indicating when the REFL DC has dipped).  That trigger is now AND-ed with a new zero crossing trigger, to make the final trigger decision.  For the zero crossing triggering, there is a small matrix (C1:LSC-ZERO_CROSS_MTRX) to choose what REFL 1f signal you'd like to use (in order, REFL11I, REFL11Q, REFL55I, REFL55Q).  The absolute value of this is compared to a threshold, which is set with the epics value C1:LSC-ZERO_CROSS_THRESH.  So, if the absolute value of your chosen RF signal is lower than the threshold, this outputs a 1, which is AND-ed by the usual schmidt trigger. 

At this moment, the input and output switches of the new filter bank are off, and the gain is set to zero.  Also, the zero crossing selection matrix is all zeros, and the threshold is set to 1e9, so it is always triggered, which means that effectively MC filter bank just has it's usual, old triggering situation.

  11006   Wed Feb 11 18:44:53 2015 manasaUpdateGeneralOptical fiber module

I have moved the optical fiber module for FOL to the PSL table. It is setup on the optical table right now for testing.

Once tests are done, the box will move to the rack inside the PSL enclosure. 

While doing any beat note alignment, please watch out for the loose fibers at the north side of the PSL enclosure until they are sheilded securely (probably tomorrow morning).

  11005   Wed Feb 11 18:11:46 2015 KojiSummaryLSC3f modulation cancellation

33MHz sidebands can be elliminated by careful choice of the modulation depths and the relative phase between the modulation signals.
If this condition is realized, the REFL33 signals will have even more immunity to the arm cavity signals because the carrier signal will lose
its counterpart to produce the signal at 33MHz.

Formulation of double phase modulation

m1: modulation depth of the f1 modulation
m2: modulation depth of the f2 (=5xf1) modulation

The electric field of the beam after the EOM

E=E_0 \exp \left[ {\rm i} \Omega t + m_1 \cos \omega t +m_2 \cos 5 \omega t \right ]
\flushleft = {\it E}_0 e^{{\rm i} \Omega t} \\ \times \left[ J_0(m_1) + J_1(m_1) e^{{\rm i} \omega t}- J_1(m_1) e^{-{\rm i} \omega t} + J_2(m_1) e^{{\rm i} 2\omega t}+ J_2(m_1) e^{-{\rm i} 2\omega t} + J_3(m_1) e^{{\rm i} 3\omega t}- J_3(m_1) e^{-{\rm i} 3\omega t} + \cdots \right] \\ \times \left[ J_0(m_2) + J_1(m_2) e^{{\rm i} 5 \omega t}- J_1(m_2) e^{-{\rm i} 5 \omega t} + \cdots \right]
\flushleft = {\it E}_0 e^{{\rm i} \Omega t} \\ \times \left\{ \cdots + \left[ J_3(m_1) J_0(m_2) + J_2(m_1) J_1(m_2) \right] e^{{\rm i} 3 \omega t} - \left[ J_3(m_1) J_0(m_2) + J_2 (m_1) J_1(m_2) \right] e^{-{\rm i} 3 \omega t} + \cdots \right\}

Therefore what we want to realize is the following "extinction" condition
J_3(m_1) J_0(m_2) + J_2(m_1) J_1(m_2) = 0

We are in the small modulation regime. i.e. J0(m) = 1, J1(m) = m/2, J2(m) = m2/8, J3(m) = m3/48
Therefore we can simplify the above exitinction condition as

m_1 + 3 m_2 = 0

m2 < 0 means the start phase of the m2 modulation needs to be 180deg off from the phase of the m1 modulation.

E = E_0 \exp\left\{ {\rm i} [\Omega t + m_1 \cos \omega t + \frac{m_1}{3} \cos (5 \omega t + \pi)] \right \}

Field amplitude m1=0.3, m2=-0.1 m1=0.2, m2=0.2
Carrier 0.975 0.980
1st order sidebands 0.148 9.9e-2
2nd 1.1e-3 4.9e-3
3rd 3.5e-7 6.6e-4
4th 7.4e-3 9.9e-3
5th 4.9e-2 9.9e-2
6th 7.4e-3 9.9e-3
7th 5.6e-4 4.9e-4
8th 1.4e-5 4.1e-5
9th 1.9e-4 5.0e-4
10th 1.2e-3 4.9e-3
11th 1.9e-4 5.0e-4
12th 1.4e-5 2.5e-5
13th 4.7e-7 1.7e-6
14th 3.1e-6 1.7e-5
15th 2.0e-5 1.6e-4

 

  11004   Wed Feb 11 18:07:42 2015 ericqUpdateLSCRFPD spectra

After some discussion with Koji, I've asked Steve to order some SBP-30+ bandpass filters as a quick and cheap way to help out REFL33. (Also some SBP-60+ for 55MHz, since we only have 1*fmod and 2*fmod bandpasses here in the lab). 

  11003   Wed Feb 11 17:31:11 2015 ericqUpdateLSCRFPD spectra

For future reference, I've taken spectra of our various RFPDs while the PRMI was sideband locked on REFL33, using a 20dB RF coupler at the RF input of the demodulator boards. The 20dB coupling loss has been added back in on the plots. Data files are attached in a zip.

Exceptions: 

  • The REFL165 trace was taken at the input of the amplifier that immediately preceeds the demod board. 
  • The 'POPBB' trace was taken with the coupler at the input of the bias tee, that leads to an amplifier, then splitter, then the 110 and 22 demod boards. 

I also completely removed the cabling for REFLDC -> CM board, since it doesn't look like we plan on using it anytime in the immediate future. 

Attachment 1: REFL.png
REFL.png
Attachment 2: AS.png
AS.png
Attachment 3: POP.png
POP.png
Attachment 4: 2015-02-PDSpectra.zip
  11002   Wed Feb 11 16:49:41 2015 ranaUpdateelogELOGD restarted

No elog response from outside and no elogd process frownon nodus, so I restarted it using 'start-elog.csh'.

  11001   Wed Feb 11 04:08:53 2015 JenneUpdateLSCNew Locking Paradigm?

[Rana, Jenne]

While meditating over what to do about the fact that we can't seem to hold PRMI lock while reducing the CARM offset, we have started to nucleate a different idea for locking

We aren't sure if perhaps there is some obvious flaw (other than it may be tricky to implement) that we're not thinking about, so we invite comments.  I'll make a cartoon and post it tomorrow, but the idea goes like this.....

Can we use ALS to hold both CARM and DARM by actuating on the ETMs, and sit at (nominally) zero offset for all degrees of freedom?  PRMI would need to be stably held with 3f signals throughout this process. 

1) Once we're close to zero offset, we should see some PDH signal in REFL11.  With appropriate triggering (REFLDC goes low, and REFL11I crosses zero), catch the zero crossing of REFL11I, and feed it back to MC2. We may want to use REFL11 normalized by the sum of the arm transmissions to some power (1, 0.5, or somewhere in between may maximize the linear range even more, according to Kiwamu).  The idea (very similar to the philosophy of CESAR) is that we're using ALS to start the stabilization, so that we can catch the REFL11 zero crossing. 

2) Now, the problem with doing the above is that actuating on the mode cleaner length will change the laser frequency.  But, we know how much we are actuating, so we can feed forward the control signal from the REFL11 carm loop to the ALS carm loop.  The goal is to change the laser frequency to lock it to the arms, without affecting the ALS lock.  This is the part where we assume we might be sleepy, and missing out on some obvious reason why this won't work.

3) Once we have CARM doubly locked (ALS pushing on ETMs, REFL11 pushing on MC/laser frequency), we can turn off the ALS system. Once we have the linear REFL11 error signal, we know that we have enough digital gain and bandwidth to hold CARM locked, and we should be able to eek out a slightly higher UGF since there won't be as many digital hops for the error signal to transverse. 

4) The next step is to turn on the high bandwidth common mode servo.  If ALS is still on at this point, it will get drowned out by the high gain CM servo, so it will be effectively off. 

5) Somewhere in here we need to transition DARM to AS55Q.  Probably that can happen after we've turned on the digital REFL11 path, but it can also probably wait until after the CM board is on.

The potential show-stoppers:

Are we double counting frequency cancellation or something somewhere?  Is it actually possible to change the laser frequency without affecting the ALS system?

Can we hold PRMI lock on 3f even at zero CARM offset?  Anecdotally from a few trials in the last hour or so, it seems like coming in from negative carm offset is more successful - we get to slightly higher arm powers before the PRMI loses lock.  We should check if we think this will work in principle and we're just saturating something somewhere, or if 3f can't hold us to zero carm offset no matter what.

A note on technique:  We should be able to get the transfer function between MC2 actuation and ALS frequency by either a direct measurement, or Wiener filtering.  We need this in order to get the frequency subtraction to work in the correct units.

  11000   Wed Feb 11 03:41:12 2015 KojiUpdateLSCPRC error signal RF spectra

As the measurements have been done under feedback control, the lower RF peak height does not necessary mean
the lower optical gain although it may be the case this time.

These non-33MHz signals are embarassingly high!
We also need to check how these non-primary RF signals may cause spourious contributions in the error signals,
including the other PDs.

  10999   Wed Feb 11 02:42:05 2015 JenneUpdateLSCPRC error signal RF spectra

Since we're having trouble keeping the PRC locked as we reduce the CARM offset, and we saw that the POP22 power is significantly lower in the 25% MICH offset case than without a MICH offset, Rana suggested having a look at the RF spectra of the REFL33 photodiode, to see what's going on. 

The Agilent is hooked up to the RF monitor on the REFL33 demod board.  The REFL33 PD has a notch at 11MHz and another at 55MHz, and a peak at 33MHz. 

We took a set of spectra with MICH at 25% offset, and another set with MICH at 15% offset.  Each of these sets has 4 traces, each at a different CARM offset.  Out at high CARM offset, the arm power vs. CARM offset is pretty much independent of MICH offset, so the CARM offsets are roughly the same between the 2 MICH offset plots. 

What we see is that for MICH offset of 25%, the REFL33 signal is getting smaller with smaller CARM offset!!  This means, as Rana mentioned earlier this evening, that there's no way we can hold the PRC locked if we reduce the CARM offset any more. 

However, for the MICH offset 15% case, the REFL 33 signal is getting bigger, which indicates that we should be able to hold the PRC.  We are still losing PRC lock, but perhaps it's back to mundane things like actuator saturation, etc. 

The moral of the story is that the 3f locking seems to not be as good with large MICH offsets.  We need a quick Mist simulation to reproduce the plots below, to make sure this all jives with what we expect from simulation.

For the plots, the blue trace has the true frequency, and each successive trace is offset in frequency by a factor of 1MHz from the last, just so that it's easier to see the individual peak heights.

Here is the plot with MICH at 25% offset:

And here is the plot with MICH at 15% offset:

Note that the analyzer was in "spectrum" mode, so the peak heights are the true rms values.  These spectra are from the monitor point, which is 1/10th the value that is actually used.  So, these peak heights (mVrms level) times 10 is what we're sending into the mixer.  These are pretty reasonable levels, and it's likely that we aren't saturating things in the PD head with these levels. 

The peaks at 100MHz, 130MHz and 170MHz that do not change height with CARM offset or MICH offset, we assume are some electronics noise, and not a true optical signal.

Also, a note to Q, the new netgpib scripts didn't write data in a format that was back-compatible with the old netgpib stuff, so Rana reverted a bunch of things in that directory back to the most recent version that was working with his plotting scripts.  sorry.

 

Attachment 1: REFL33_25.pdf
REFL33_25.pdf
Attachment 2: REFL33_15.pdf
REFL33_15.pdf
  10998   Wed Feb 11 00:07:54 2015 ranaUpdateLSCLock Loss plot

Here is a lock loss from around 11 PM tonight. Might be due to poor PRC signals.  \oint {\frac{\partial PRCL}{\partial x}}

This is with arm powers of ~6-10. You can see that with such a large MICH offset, POP22 signal has gone done to zero. Perhaps this is why the optical gain for PRCL has also dropped by a factor of 30 crying.

This seems untenable no. We must try this whole thing with less MICH offset so that we can have a reasonable PRCL signal.cool

Attachment 1: 1107673198.png
1107673198.png
  10997   Tue Feb 10 18:37:17 2015 ericqUpdateASCQPD ASC ready to go

I've remeasured the QPD ASC sensing coefficients, and figured out what I did weird with the actuation coefficients. I've rearranged the controller filters to be able to turn on the boost in a triggered way, and written Up/Down scripts that I've tested numerous times, and Jenne has used as well; they are exposed on the ASC screen. 

All four loops (2 arms * pit,yaw), have their gains set for 8Hz UGF, and have 60 degrees of phase margin. The loop shape is the same as the previous ELOG post. Here is the current on/off performance. The PDH signals (not shown, but in attached xml) show no extra noise, and the low frequency RIN goes down a bit, whic is good. The oplevs error signals are a bit noisier, but I suppose that's unavoidable. The Y-arm performs a bit better than the X-arm. 

The up/down scripts don't touch the filters' trigger settings at all, just handles switching the input and output and clearing history. FM1 contains the boost which is intended to have a longer trigger delay than the filters themselves.

Attachment 1: Feb10_loops_offOn.png
Feb10_loops_offOn.png
Attachment 2: Feb10_newLoops_offOn.xml.zip
  10996   Tue Feb 10 16:01:21 2015 manasaUpdateGeneralAUX X fiber coupled 72%

Plan C finally worked. We have 1.454mW of AUX X light at the PSL table (2mW incident on the fiber coupler). 

Attached is the layout of the telescope.

What I did:

I stuck in Lens 1 (f=200mm) and measured the beam width after the focus of the lens at several points. I fit the data and calculated the beam waist and its position after this lens. 

I used the calculated waist and matched it with an appropriate lens and target (fiber coupler) distance. I calculated the maximum coupling efficiency to be 77% for Lens 2 with f=50mm and the fiber coupler placed at 20cm from the waist of Lens1. I was able to obtain 72% coupling after putting the telescope together.

I locked the arms, ran ASS and brought back GTRX to its usual optimum value of ~0.5 counts after closing. We also have the X arm beatnote on the spectrum analyzer.

Notes:

There are still a couple of things to fix. The rejected beam from the beam sampler has to be dumped using a razor blade.

 

  10995   Tue Feb 10 13:48:58 2015 manasaUpdateLSCProbable cause for headaches last night

I found the PSL enclosure open (about a feet wide) on the north side this morning. I am assuming that whoever did the X beatnote alignment last night forgot to close the door to the enclosure before locking attempts frown

Quote:

Unfortunately, we only had one good CARM offset reduction to powers of about 25, but then my QPD loop blew it. We spent the vast majority of the night dealing with headaches and annoyances. 

Things that were a pain:

  • If TRX is showing large excursions after finding resonance, there is no hope. These translate into large impulses while reducing the CARM offset, which the PRMI has no chance of handling. The first time aligning the green beat did not help this. For some reason, the second time did, though the beatnote amplitude wasn't increased noticibly. 
    • NOTICE: We should re-align the X green beatnote every night, after a solid ASS run, before any serious locking work. 
    • Afterwards, phase tracker UGFs (which depend on beatnote amplitude, and thereby frequency) should be frequently checked. 
  • We suffered some amount from ETMX wandering. Not only for realigning between lock attempts, but on one occasion, with CARM held off, GTRX wandered to half its nominal value, leading to a huge effective DARM offset, which made it impossible to lock MICH with any reasonble power in the arms. Other times, simply turning off POX/POY locking, after setting up the beatnotes, was enough to significantly change the alignment. 
  • IMC was mildly tempermental, at its worst refusing to lock for ~20min. One suspicion I have is that when the PMC PZT is nearing its rail, things go bad. The PZT voltage was above 200 when this was happening, after relocking the PMC to ~150, it seems ok. I thing I've also had this problem at PZT voltages of ~50. Something to look out for. 

Other stuff:

  • We are excited for the prospect of the FOL system, as chasing the FSS temperature around is no fun. 
  • UGF servo triggering greatly helps the PRMI reacquire if it briefly flashes out, since the multipliers don't run away. This exacerbated the ALS excursion problem. 
  • Using POPDC whitening made it very tough to hold the PRMI. Maybe because we didn't reset the dark offset...?

 

  10994   Tue Feb 10 03:09:02 2015 ericqUpdateLSCSome locking thoughts on PRMI

Unfortunately, we only had one good CARM offset reduction to powers of about 25, but then my QPD loop blew it. We spent the vast majority of the night dealing with headaches and annoyances. 

Things that were a pain:

  • If TRX is showing large excursions after finding resonance, there is no hope. These translate into large impulses while reducing the CARM offset, which the PRMI has no chance of handling. The first time aligning the green beat did not help this. For some reason, the second time did, though the beatnote amplitude wasn't increased noticibly. 
    • NOTICE: We should re-align the X green beatnote every night, after a solid ASS run, before any serious locking work. 
    • Afterwards, phase tracker UGFs (which depend on beatnote amplitude, and thereby frequency) should be frequently checked. 
  • We suffered some amount from ETMX wandering. Not only for realigning between lock attempts, but on one occasion, with CARM held off, GTRX wandered to half its nominal value, leading to a huge effective DARM offset, which made it impossible to lock MICH with any reasonble power in the arms. Other times, simply turning off POX/POY locking, after setting up the beatnotes, was enough to significantly change the alignment. 
  • IMC was mildly tempermental, at its worst refusing to lock for ~20min. One suspicion I have is that when the PMC PZT is nearing its rail, things go bad. The PZT voltage was above 200 when this was happening, after relocking the PMC to ~150, it seems ok. I thing I've also had this problem at PZT voltages of ~50. Something to look out for. 

Other stuff:

  • We are excited for the prospect of the FOL system, as chasing the FSS temperature around is no fun. 
  • UGF servo triggering greatly helps the PRMI reacquire if it briefly flashes out, since the multipliers don't run away. This exacerbated the ALS excursion problem. 
  • Using POPDC whitening made it very tough to hold the PRMI. Maybe because we didn't reset the dark offset...?
  10993   Tue Feb 10 02:55:29 2015 JenneConfigurationModern ControlQuacky filters

I discovered that somehow my Wiener filters that show up in Foton are not the same as what I have in Matlab-land. 

I have plotted each of my 3 filters that I'm working with right now (T-240 X, Y and Z for PRCL Pitch) at several stages in the filter creation process.  Each plot has:

  • Blue trace is the Wiener filter that I want, after the vectfit process.
  • Green trace is the frequency response of the SOS filters that are created by autoquack (really, quack_to_rule_them_all, which is called by autoquack).  The only thing that happens in matlab after this is formatting the coefficients into a string for writing to foton.
  • Red trace is the exported text file from foton.

It's not just a DC gain issue - there's a frequency dependent difference between these filters.  Why?? 

It's not frequency warping from changing between analog and digital filters.  The sample rate for the OAF model is 2048Hz, so the effect is small down at low frequencies.  Also, the green trace is already discretized, so if it were freq warping, we'd see it in the green as well as red, which clearly we don't.

Has anyone seen this before?  I'm emailing my seismic friends who deal in quack more than I do (BLantz and JKissel, in particular) to see if they have any thoughts.

Also, while I'm asking questions, can autoquack clear filters?  Right now I'm overwriting old filters with zpk([],[],1)'s, which isn't quite the same.  (I need this because the Wiener code needs more than one filter module to fit all of the poles I need, and it decides for itself how many FMs it needs by comparing the length of the poles vector with 20.  If one iteration needs 4 filter modules, but the next iteration only wants 3, I don't want to leave in a bogus 4th filter. 

Here are the plots:

(The giant peak at ~35Hz in the Z-axis fiilter is what tipped me off that something funny was going on)

  10992   Tue Feb 10 02:40:54 2015 JenneUpdateLSCSome locking thoughts on PRMI

[EricQ, Jenne]

We wanted to make the PRMI lock more stable tonight, which would hopefully allow us to hold lock much longer.  Some success, but nothing out-of-this-world.

We realized late last week that we haven't been using the whitening for the ASDC and POPDC signals, which are combined to make the MICH error signal.  ASDC whitening is on, and seems great.  POPDC whitening (even if turned on after lock is acquired) seems to make the PRMI lock more fussy.  I need to look at this tomorrow, to see if we're saturating anything when the whitening is engaged for POPDC.

One of the annoying things about losing the PRMI lock (when CARM and DARM have kept ALS lock) is that the UGF servos wander off, so you can't just reacquire the lock.  I have added triggering to the UGF servo input, so that if the cavity is unlocked (really, untriggered), the UGF servo input gets a zero, and so isn't integrating up to infinity.  It might need a brief "wait" in there, since any flashes allow signal through, and those can add up over time if it takes a while for the PRMI to relock.  UGF screens reflect this new change.

ELOG V3.1.3-